What happens if we graph both f and f^{-1} on the same set of axes, using the x-axis for the input to both f and f^{-1} ?

[Suggestion: go to www.desmos.com/calculator and type y=x^3 {-2 < x < 2}, y=x^{1/3} {–2 < x < 2}, and y = x {–2 < x < 2}, and describe the relationship between the three curves.] Then post your own example discussing the difficulty of graph both f and f^{-1} on the same set of axes.

Suppose f:R \rightarrow R is a function from the set of real numbers to the same set with f(x)=x+1. We write f^{2} to represent f \circ f and f^{n+1}=f^n \circ f. Is it true that f^2 \circ f = f \circ f^2? Why? Is the set {g:R \rightarrow R l g \circ f=f \circ g} infinite? Why?
1
Expert's answer
2020-12-14T16:11:18-0500

a) What happens if we graph both f and "f^{-1}" on the same set of axes, using the x-axis for the input to both f and "f^{-1}"?


first of all f and "f^{-1}" are inverses of each other.

Let's see if we graph both f and "f^{-1}" on the same set of axes. Let's denote "f^{-1}=g"

  • (f o g)(x) = x , x in the domain of g and (g o f)(x) = x , x in the domain of f
  • the domain of f is equal to the range of g and the range of f is equal to the domain of g.
  • their graphs are symmetrical reflections of each other with respect to the line y = x.

b ) [Suggestion: go to www.desmos.com/calculator and type y="x^3"{-2 < x < 2}, y="x^{1\/3}" {–2 < x < 2}, and y = x {–2 < x < 2}, and describe the relationship between the three curves.] Then post your own example discussing the difficulty of graph both f and "f^{-1}" on the same set of axes.


y="x^3" {-2 < x < 2}, y="x^{1\/3}" {–2 < x < 2}, and y = x {–2 < x < 2}. Let's see their graphs on www.desmos.com/calculator.





The following properties can be seen clearly from the graph

  • let g = "x^3" and g ="x^{1\/3}"

We find f(g(x)) = "(x^{1\/3})^3=x" and g(f(x)) = "(x^3)^{1\/3}" = x

f(g(x)) = g(f(x))

  • let's find their domain and the range:

the domain of "x^3" is (-"\infty", +"\infty" ) and its range is (-"\infty" , +"\infty" ),

the domain of "x^{1\/3}" is (-"\infty" , +"\infty" ) and its range is (-"\infty" , +"\infty" )

it can be seen the domain of f is equal to the range of g and the range of f is equal to the domain of g.

  • the functions are a reflection of each other with respect to the graph of y = x.

my own example

y ="x^2" and y = "\sqrt{x}" , y = - "\sqrt{x}"





  • let g = "x^2" and g ="x^{1\/2}" and "-x^{1\/2}"

we find f(g(x)) = "(x^{1\/2})^2=x" and g(f(x)) = "(x^2)^{1\/2}" = x

f(g(x)) = g(f(x))

  • lets find their domain and range

the domain of "x^2" is (-"\infty" , +"\infty" ) and the range is [0, +"\infty" )

"-x^{1\/2}" -x^(1/2) and "x^{1\/2}" have the domain [0, +"\infty" ) and the range (-"\infty" , +"\infty" )

it can be seen the domain of f is equal to the range of g and the range of f is equal to the domain of g

  • the functions are a reflection of each other with respect to the graph of y = x.

c) Suppose f:R "\rightarrow" R is a function from the set of real numbers to the same set with f(x)=x+1. We write "f^{2}" to represent f "\circ" f and "f^{n+1}=f^n" "\circ" f.

Is it true that "f^2" "\circ" f = f"\circ" "f^2"? Why? Is the set {g:R "\rightarrow" R l g"\circ" f=f o g} infinite? Why?

"f^2" = f "\circ" f = x+1+1 = x+2

"f^2 \circ f" = x+1+2 = x+3

"f \circ f^2" = x+2+1 = x+3, so yes they are the same;

"f^3" ="f^2 \circ f" = x+3

"f^4=f^3 \circ f" = x+4

...

Suppose "f^n" = x+n,

"f^{n+1}=f^n \circ f" = x+1+n = x+(n+1)

so the pattern holds by induction,

F "\circ" Finv = x so when you graph them on

the same plot, they shall be SYMMETRIC about

the line y=x